The cost price, ₦2,500 of a radio is reduced by ₦750.
a) Find the ratio by which the price is reduced.
b) Express the price reduction as a rate of kobo in naira.

Answer by Theo(13342)

regular price is 2500
discount is 750
sale price is 2500 - 750 = 1750.
the ratio of the reduction in price to the original price is 750/2500 = .3
that's because 2500 - .3 * 2500 = 1750
if you factor out the 2500, then you get 2500 * (1 - .3) = 2500 * .7 = 1750.

from britannica, i get:
The naira is divided into 100 kobo. The naira was introduced in 1973, when the country decimalized its monetary system and substituted the naira for the Nigerian pound (the country used the British pound sterling when it was a British colony), which was divided into shillings.

that's the official rate as far as i can tell.

if the reduction is in naira, then a reduction of 750 naira is equivalent to a reduction of 75,000 kobo, because 750 * 100 = 75,000.